# Causality of higher-spin interactions on the (A)dS lightcone, with application to the static patch

### v1

@@ Appendix B Comparing coupling constants between the Fronsdal and lightcone formalisms
In this Appendix, we derive the scaling constant $a$ in the proportionality (28) between the chiral cubic couplings of different helicities. We do this for the case of type-A HS gravity, by comparing to its holographic dual – the free $O(N)$ vector model. It is sufficient to consider the case $(h_{1},h_{2},h_{3})=(0,0,\pm s)$, and to focus on the highest-derivative terms (equivalently, the flat limit, or the large-$z$ limit). In this sector, the lightcone Lagrangian to cubic order is given by (9),(22)-(24),(28)-(32) as:
@@
$\begin{split}\mathcal{L}={}&\frac{1}{2}\Phi_{0}\Box\Phi_{0}+\Phi_{-s}\Box\Phi_{s}\\ &+\frac{Ca^{s}z^{s-1}}{(s-1)!}\,\Phi_{0}\left(\big(q^{\alpha_{1}}\partial_{\alpha_{1}\dot{\alpha}_{1}}\dots q^{\alpha_{s}}\partial_{\alpha_{s}\dot{\alpha}_{s}}\Phi_{0}\big)\left(\frac{q^{\beta_{1}}\partial_{\beta_{1}}{}^{\dot{\alpha}_{1}}\dots q^{\beta_{s}}\partial_{\beta_{s}}{}^{\dot{\alpha}_{s}}}{(\ell\cdot\partial)^{s}}\,\Phi_{s}\right)+c.c.\right)\ ,\end{split}$ (145)
@@
with $C$ an overall coupling constant. Here, we used the flat limit (i.e. neglected terms where the derivatives act on the $z$ factor) to replace $\mathbb{P}\to p^{(2)}_{\dot{\alpha}}p^{(3)\dot{\alpha}}$ via integration-by-parts, making the first $\Phi_{0}$ factor derivative-free. We also used the fact that in the type-A theory, $C$ is real Skvortsov:2018uru .
@@
A covariant analog of the Lagrangian (145) was worked out in Bekaert:2015tva (and for all spins $s_{1},s_{2},s_{3}$ in Sleight:2016dba ), using the holographic correlators to fix the cubic couplings. The field variables in Bekaert:2015tva are a scalar $\varphi$ and a spin-$s$ Fronsdal field $\varphi_{\mu_{1}\dots\mu_{s}}$ Fronsdal:1978vb . We take $\varphi_{\mu_{1}\dots\mu_{s}}$ to be transverse and traceless. Since in this paper we raise/lower indices with the flat metric $\eta_{\mu\nu}$, we will be careful to define $\varphi_{\mu_{1}\dots\mu_{s}}$ with lower indices, and to include explicit powers of $g^{\mu\nu}=z^{2}\eta^{\mu\nu}$ when its indices are raised. We also include in the Lagrangian an explicit factor of $\sqrt{-g}=1/z^{4}$. On the other hand, we use the flat limit to write AdS covariant derivatives simply as partial derivatives $\partial_{\mu}$. With these settings, the cubic Lagrangian of Bekaert:2015tva ; Sleight:2016dba reads:
@@
$\begin{split}\mathcal{L}={}&\frac{1}{2z^{2}}\Big(\varphi\Box\varphi+z^{2s}\eta^{\mu_{1}\nu_{1}}\dots\eta^{\mu_{s}\nu_{s}}\varphi_{\mu_{1}\dots\mu_{s}}\Box\varphi_{\nu_{1}\dots\nu_{s}}\Big)\\ &+\frac{2^{s/2}C^{\prime}z^{2s-4}}{(s-1)!}\,\eta^{\mu_{1}\nu 1}\dots\eta^{\mu_{s}\nu_{s}}\varphi\,\varphi_{\mu_{1}\dots\mu_{s}}\partial_{\nu_{1}}\dots\partial_{\nu_{s}}\varphi\ ,\end{split}$ (146)
@@
In the flat limit (i.e. taking all derivatives as flat), it’s easy to see that the ansatz (148) is transverse and traceless. Plugging the ansatz (147)-(148) into the kinetic terms of the convariant Lagrangian (146), we recover the kinetic terms of the lightcone Lagrangian (145). Next, we plug (147)-(148) into the cubic vertex of (146). The result matches with (145), if we identify:

### v2

@@ Appendix B Comparing coupling constants between the Fronsdal and lightcone formalisms
In this Appendix, we derive the scaling constant $a$ in the proportionality (28) between the chiral cubic couplings of different helicities. We do this for the case of type-A HS gravity, by comparing to its holographic dual – the free $O(N)$ vector model. It is sufficient to consider vertices of the form $(h_{1},h_{2},h_{3})=(0,0,\pm s)$, and to focus on the highest-derivative terms (equivalently, the flat limit, or the large-$z$ limit).
Let us start by writing the most general real lightcone Lagrangian to cubic order, with its chiral couplings subject to the pattern (28) derived from the chiral sector’s quartic consistency. Using (9),(22)-(24),(28)-(32), this can be written as:
$\mathcal{L}={}$ $\frac{1}{2}\Phi_{0}\Box\Phi_{0}+\sum_{s>0}\Phi_{-s}\Box\Phi_{s}$ (145)
@@
$+\sum_{s>0}\frac{z^{s-1}}{(s-1)!}\,\Phi_{0}\left(Ca^{s}\big(q^{\alpha_{1}}\partial_{\alpha_{1}\dot{\alpha}_{1}}\dots q^{\alpha_{s}}\partial_{\alpha_{s}\dot{\alpha}_{s}}\Phi_{0}\big)\left(\frac{q^{\beta_{1}}\partial_{\beta_{1}}{}^{\dot{\alpha}_{1}}\dots q^{\beta_{s}}\partial_{\beta_{s}}{}^{\dot{\alpha}_{s}}}{(\ell\cdot\partial)^{s}}\,\Phi_{s}\right)+c.c.\right)\ ,$
@@
where $C$ and $a$ are (for the moment) complex constants, and we display only the vertices with $(h_{1},h_{2},h_{3})=(0,0,\pm s)$. The fields $\Phi_{h}$ ($h=0,\pm s$) satisfy the reality conditions $\Phi_{h}^{\dagger}=\Phi_{-h}$. When writing the vertices in (145), we used the flat limit (i.e. neglected terms where the derivatives act on the $z$ factor) to replace $\mathbb{P}\to p^{(2)}_{\dot{\alpha}}p^{(3)\dot{\alpha}}$ via integration-by-parts, making the first $\Phi_{0}$ factor derivative-free.
Now, in the type-A theory, the overall chiral coupling $C$ is real Skvortsov:2018uru , and can always be made positive by flipping the sign of all the fields $\Phi_{h}\to-\Phi_{h}$. Furthermore, as discussed in the main text, we can use phase rescalings $\Phi_{h}\to e^{ih\theta}\Phi_{h}$ to make $a$ real and positive as well. We can then take $C$ and $a$ outside the “+c.c.{}+c.c.” parentheses. Writing out both chiral and anti-chiral vertices explicitly, we obtain:
$\mathcal{L}={}$ $\frac{1}{2}\Phi_{0}\Box\Phi_{0}+\sum_{s>0}\Phi_{-s}\Box\Phi_{s}$ (146)
$+C\sum_{s>0}\frac{a^{s}z^{s-1}}{(s-1)!}\,\Phi_{0}\left(\big(q^{\alpha_{1}}\partial_{\alpha_{1}\dot{\alpha}_{1}}\dots q^{\alpha_{s}}\partial_{\alpha_{s}\dot{\alpha}_{s}}\Phi_{0}\big)\left(\frac{q^{\beta_{1}}\partial_{\beta_{1}}{}^{\dot{\alpha}_{1}}\dots q^{\beta_{s}}\partial_{\beta_{s}}{}^{\dot{\alpha}_{s}}}{(\ell\cdot\partial)^{s}}\,\Phi_{s}\right)\right$.
$\qquad\qquad\qquad\qquad\left.{}+\big(\bar{q}^{\dot{\alpha}_{1}}\partial_{\alpha_{1}\dot{\alpha}_{1}}\dots\bar{q}^{\dot{\alpha}_{s}}\partial_{\alpha_{s}\dot{\alpha}_{s}}\Phi_{0}\big)\ \left(\frac{\bar{q}^{\dot{\beta}_{1}}\partial^{\alpha_{1}}{}_{\dot{\beta}_{1}}\dots\bar{q}^{\dot{\beta}_{s}}\partial^{\alpha_{s}}{}_{\dot{\beta}_{s}}}{(\ell\cdot\partial)^{s}}\,\Phi_{-s}\right)\right)\ .$
Before comparing to the holographic dual, let us emphasize the conceptual status of the Lagrangian (146). First, note that the real Lagrangian (146) can be considered as-is, irrespective of its origin from the chiral theory. This makes it clear that different positive values of $a$ are indeed distinct: there is no further rescaling of the fields that can change one Lagrangian (146) into another with a different positive $a$. For each of these Lagrangians, the chiral sector’s quartic consistency conditions in the flat limit are satisfied. On the other hand, in the holographic CFT dual, there is no 1-parameter freedom that would correspond to different choices of $a$. This implies that quartic consistency does, after all, require a particular value of $a$, but that this stricter consistency condition cannot be accessed from just the flat limit: the full AdS consistency analysis would probably need to be carried out.
@@
With this understanding, we are ready to read off the correct value of $a$ from the simplest holographic dual – the free $O(N)$ vector model. To perform the comparison, we will need the covariant version of the Lagrangian (146). This was worked out in Bekaert:2015tva (and for all spins $s_{1},s_{2},s_{3}$ in Sleight:2016dba ), using the holographic correlators to fix the cubic couplings. The field variables in Bekaert:2015tva are a scalar $\varphi$ and spin-$s$ Fronsdal fields $\varphi_{\mu_{1}\dots\mu_{s}}$ Fronsdal:1978vb , which we take to be transverse and traceless. Since in this paper we raise/lower indices with the flat metric $\eta_{\mu\nu}$, we will be careful to define $\varphi_{\mu_{1}\dots\mu_{s}}$ with lower indices, and to include explicit powers of $g^{\mu\nu}=z^{2}\eta^{\mu\nu}$ when its indices are raised. We also include in the Lagrangian an explicit factor of $\sqrt{-g}=1/z^{4}$. On the other hand, we use the flat limit to write AdS covariant derivatives simply as partial derivatives $\partial_{\mu}$. With these settings, the cubic Lagrangian of Bekaert:2015tva ; Sleight:2016dba reads:
@@
$\begin{split}\mathcal{L}={}&\frac{1}{2z^{2}}\Big(\varphi\Box\varphi+\sum_{s>0}z^{2s}\eta^{\mu_{1}\nu_{1}}\dots\eta^{\mu_{s}\nu_{s}}\varphi_{\mu_{1}\dots\mu_{s}}\Box\varphi_{\nu_{1}\dots\nu_{s}}\Big)\\ &+C^{\prime}\sum_{s>0}\frac{2^{s/2}z^{2s-4}}{(s-1)!}\,\eta^{\mu_{1}\nu 1}\dots\eta^{\mu_{s}\nu_{s}}\varphi\,\varphi_{\mu_{1}\dots\mu_{s}}\partial_{\nu_{1}}\dots\partial_{\nu_{s}}\varphi\ ,\end{split}$ (147)
@@
In the flat limit (i.e. taking all derivatives as flat), it’s easy to see that the ansatz (149) is transverse and traceless. Plugging the ansatz (148)-(149) into the kinetic terms of the covariant Lagrangian (147), we recover the kinetic terms of the lightcone Lagrangian (146). Next, we plug (148)-(149) into the cubic vertex of (147). The result matches with (146), if we identify: